A merchant wishes to mix candy worth $5 per lb with 40 lb of candy worth $2 per lb to get a mixture that can be sold for $3 per lb. How many pounds of $5 candy should be used?
Answer by ptaylor(2198)


Let x= amount of $5 per lb candy
Then 40-x=amount of $2 per lb candy
Now we know that the value of $5 candy (5x) plus the value of the $2 candy (2(40-x)) equals the value of the final mixture $3(40). So our equation to solve is:
5x+2(40-x)=3(40) get rid of parens
5x+80-2x=120 subtract 80 from both sides
5x+80-80-2x=120-80 collect like terms
3x=40 divide both sides by 3
x=13.3333 lb----------------------------------amount of $5 per lb candy
40-x=40-13.3333=26.6667 lb----------------------amount of $2 candy
Check
13.3333($5)+26.6667($2)=$3(40)
66.6665+53.3333=120
~120=120
Hope this helps----ptaylor